OpenStudy (anonymous):

If k is a constant such that \(x^{3}+4x^{2}+kx-12\) is divisible by x+3, then k=?

hartnn (hartnn):

just plug in x =-3 and equate it to 0

hartnn (hartnn):

if x+3 is a factor then, x =-3 is a root

OpenStudy (anonymous):

i got 1 but it's not correct :O

hartnn (hartnn):

(-3)^3+4^2-3k−12 = 0

hartnn (hartnn):

what is that 4^2 doing there

hartnn (hartnn):

can you please verify your question again /

hartnn (hartnn):

is it x^2 and not 4^2

hartnn (hartnn):

if it is x^2 i got k =-10

OpenStudy (anonymous):

4x^2

hartnn (hartnn):

\(x^3+4x^2 +kx -12 =0, for \: x = -3\) i got x=-1 too

hartnn (hartnn):

k =-1

hartnn (hartnn):

how did you get +1 ?

OpenStudy (anonymous):

:O -3+k(-3)=0 .... oh i got it :O

hartnn (hartnn):

excellent! :)

OpenStudy (anonymous):

thanks :D